OpenStudy (anonymous):

Mandy's Candies is most famous for their homemade fudge and homemade toffee. They make an average of 150 pounds of fudge and toffee combined each month. They sell a pound of fudge for $8.54 and a pound of toffee for $7.97. Mandy's averages $1,232.55 each month in fudge and toffee sales. On average, how many pounds each of fudge and toffee does Mandy's Candies sell in one month?

OpenStudy (anonymous):

Let pounds of fudge = f and pounds of toffee = t. 1 pound of fudge is $8.54 and 1 pound of toffee is $7.97. The total sales are $1232.55. The following are true. (1) f+t=150, which is saying that the pounds of f and pounds of t together are 150 pounds. (2) $8.54*f +$7.97*t=$1232.55, which is saying that the pounds of fudge we have times the cost for 1 pound of fudge plus the same for toffee equals $1232.55. Now we have a system of equations: f+t=150 8.54f+7.97t=1232.55